The potential energy of a 0.2 kg particle moving along the x axis is given by:


U(x) = 8x2 - 2x4


where U is in Joules and x in meters. When the particle is at x = 10 m, what will be its acceleration?

Grade:12

1 Answers

vikas askiitian expert
13 years ago

F = -du/dx 

   ={8x^3 - 16x}

 F=ma

 a=40(x^3 - 2x)

at x=10

 a=3.92*10^4